# Comparing Inferential Strategies of Humans and Large Language Models in Deductive Reasoning

### v1

@@ 1 Introduction
Models are often right but for the wrong reasons: the accuracy of a model, that is the number of correct final answers, does not reflect whether its reasoning is sound, i.e. logically follows from the statements at hand.
@@
In contrast to human reasoners, models occasionally adopt a symbolic strategy, where formal logical calculus is employed to solve the problems of propositional logic.
@@ 2 Strategies in Propositional Reasoning
To gain insights into the inferential processes humans employ in propositional reasoning, Van der Henst et al. (2002) conducted a series of experiments that study the behavior of participants during propositional reasoning. They formulated straightforward propositional logic problems with neutral content (the presence or absence of colored marbles in a box, similar to the problem illustrated in Figure 1) and requested participants to articulate their thought processes while engaging with these problems. Participants were permitted the use of paper and pencil for their workings. Both their verbal explanations and written responses were meticulously recorded, transcribed and analyzed thereafter. Van der Henst et al. (2002) discovered five strategies reasoners commonly utilize to navigate the problems, offering insights into their inferential mechanisms employed in propositional reasoning. In the following, we give a short description of each strategy (illustrated in Figure 2). For more details and additional examples, we refer to the original study by Van der Henst et al. (2002).
@@
Incremental Diagram. This strategy involves the creation of a comprehensive diagram that keeps track of all potential outcomes compatible with the statements of the problem. During the reasoning process, individuals progressively increment their diagrams to incorporate new information derived (see left box in Figure 2). The result is a single diagram that records a variety of possibilities compatible with the premises, often including even those that might be irrelevant to the task.111In contrast to Van der Henst et al. (2002), we in fact observe no single occurrence of the incremental diagram strategy in LLMs, despite the authors finding this strategy being employed most frequently by humans. We belief that this discrepancy stems from the use of pen and paper in human assessments, implicitly encouraging diagrammatic reasoning.
@@
Supposition Following. Reasoners employing this strategy start with a supposition, e.g. by assuming a marble of a certain color. They then trace the consequences of that supposition following from the statements at hand, as illustrated in the second box from the left of Figure 2. The result is a sequence of literals (in this case, marbles of a certain color) without logical connectives. The success of supposition following strongly depends on the supposition made by the reasoner. Some suppositions might lead to inferences that are relevant to the problem, while others might lead to irrelevant conclusions.
@@
Chain Construction. When employing this strategy, reasoners construct a chain of conditional statements derived from the premises at hand or from intermediate deductions. An example of chain construction is displayed in the central box of Figure 2. Premises are converted into a chain of conditional statements that are linked by their entities. A distinctive feature of this chain is the interconnection between conditionals, where the consequent of one conditional is the antecedent of the following.
@@
Compound Strategy. Reasoners following the compound strategy combine two or more statements to derive a new compound conclusion. This process yields a series of novel conclusions, each building upon the preceding ones. An example illustrating this strategy is given in the upper second box from the right of Figure 2. Based on the first two premises, the compound conclusion: “If red then not blue.” is inferred, and then used to draw another compound conclusion (“If red or pink then not blue.”) together with the last premise of the problem statement.
@@ 3 Experimental Setup
Task Overview. Our task setup aligns with the experiment conducted by Van der Henst et al. (2002) to allow for a fair comparison between the inferential strategies found in humans and those identified in LLMs.222More specifically, experiment one of Van der Henst et al. (2002). In particular, we evaluate each model on the 12 problems of propositional logic suggested by Van der Henst et al. (2002) (an overview of each problem can be found in Figure 5 in the appendix). In each problem, models are presented with a set of statements (or premises) and must determine whether a given conclusion logically follows (for an example, see Figure 1). Eight out of 12 problems involve three premises and a conclusion, while the remaining four problems consist of four premises leading to a conclusion. All premises, as well as the conclusions resemble either biconditionals, exclusive disjunctions or conditionals. Two problems (4 and 6) include a redundant first premise. All premises are stated such that two subsequent statements contain one proposition in common, except of two problems (11 and 12), which are arranged in a non-sequential manner. In half of the problems, the conclusions logically follow from the premises, whereas in the other half, they do not. To avoid the influence of external knowledge and ensure content neutrality, Van der Henst et al. (2002) framed the problems around the presence of colored marbles in a box, with colors assigned randomly to each entity within a problem.
@@
Evaluation Setup. We prompt each model with a system message providing context about the task they are about to solve and the format in which they should answer (for the full prompt, see Figure 5 in the appendix). Analogous to Van der Henst et al. (2002), we inform the model of its participation in an experiment designed to explore reasoning processes, and instruct it to “think aloud” as it tackles the problem. In addition to the system message, we provide a user prompt that contains the problem description. In cases where the model does not accept system messages (such as Mistral-7B-Instruct-v0.2), we prepend the content of the system message to the user prompt. To prevent biasing the model towards a certain strategy, we refrain from providing few-shot examples, as done also by Leidinger et al. (2023). Instead, we elicit reasoning through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2023). Answers for each model are generated with nucleus sampling using Llama-2-chat-hf’s default values (, temperature ), as we found this configuration to work well for all models. To account for the statistical nature of language models, we ask each model to solve the set of propositional problems across 5 random seeds, resulting in a total of 60 responses per model.
We record all answers and manually evaluate them (a total of 300 responses) for strategies employed in their reasoning (see Figure 3 for an example). For each model response, we qualitatively evaluate for strategy and soundness. That is, we manually label the inferential strategies identified, and the logical validity of the model’s reasoning. In addition, we record whether the final answer is correct. In cases of faulty reasoning, we categorize the type of error. This comprehensive manual evaluation of model responses is independently conducted by two hired students with expertise in manual data annotation. To gauge the quality of the annotations, we report an overall Cohen’s Kappa value of . For details on the inter-annotator agreement of each label, we refer to Table 2 in the appendix. Further annotated examples can be found in Appendix C. Following the recommendations put forward by Leidinger et al. (2023), we make all input prompts, model responses, code and manual annotations publicly available upon acceptance of this paper.
Model |
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |
LLaMA-2-70B |
Human Reasoner |
@@ 4 Results and Analysis
In this section, we present the results of our evaluation. We begin with a quantitative analysis of the inferential strategies employed by LLMs, as well as the logical validity of their reasoning. This is followed by a qualitative analysis providing a more in-depth examination of the models’ reasoning behavior.
@@ 4.1 Quantitative Analysis
Table 1 provides an overview of the frequencies with which large language models employ inferential strategies when navigating the problems of propositional logic described in Section 3. Our evaluation reveals that all models display strategies akin to those observed by Van der Henst et al. (2002). In particular, we find that, similar to humans, models commonly employ supposition following, chain construction and the compound strategy. In addition, we observe that models occasionally utilize the symbolic strategy, employing techniques from logical calculus to solve the tasks (see Section 2). Note that, similar to humans, models might switch from one strategy to another during a single problem, demonstrating multiple strategies within their responses (see Figure 19 in the appendix for an example). Surprisingly, we observe that distinct model families favor different inferential strategies. For instance, Zephyr-7B- predominantly employs supposition following, while Mistral-7B-Instruct is inclined towards drawing compound conclusions. In comparison, models from the Llama 2 series tend to rely on supposition following and chain construction, with negligible use of the compound strategy. Our analysis further reveals a discrepancy between the correctness of the models’ final answers and the logical soundness of their reasoning. While all models achieve an answer accuracy that approximately coincides with chance in our experimental setup, an analysis of their reasoning validity reveals a different picture: LLaMA-2-70B outperforms the other models by reasoning correctly in about 31.7% of cases, while Zephyr-7B- and Mistral-7B-Instruct produce sound reasoning in 25% of problems. We note that all models perform rather poorly on the propositional tasks, with LLaMA-2-7B failing entirely to construct sound arguments.
@@
Sound Reasoning. As previously highlighted, the accuracy of a model’s final answer does not necessarily serve as a reliable indicator of its reasoning capability. In particular, we observe that models often arrive at correct answers, but through flawed reasoning processes (refer to Figure 10 in the appendix for an illustration). Interestingly, we also find instances where models provide incorrect final answers despite reasoning correctly (for an example, see Figure 16 in the appendix). Our analysis reveals only a moderate positive correlation between the accuracy of the models’ final answers and the logical soundness of their reasoning, with a Pearson correlation coefficient and a statistically significant p-value of less than . This observation underscores the need for more nuanced evaluation procedures, particularly in multiple-choice settings, where models might select the correct answer by chance rather than through rigorous reasoning.
@@ 7 Limitations
While our work contributes to the understanding of reasoning processes in large language models by demonstrating that these models, similar to humans, employ inferential strategies in propositional logic, it encompasses several limitations that could be addressed in future work.
Task setup. Our study is constrained by a limited set of problems, designed within a fixed framework that revolves around hypothesis validation based on 3-4 statements of propositional logic. We employ a constant and neutral content, disregarding potential content-effects on the models’ reasoning behavior, as shown by Dasgupta et al. (2023). Similarly, we have not yet examined factors such as the complexity of the problems, the differences between hypothesis validation and generation, and the impact of logical connectives utilized in the premises. We belief that these factors are worth investigating and leave a detailed examination to future work.
Evaluation Framework. The extent of our manual evaluation is limited by both the number of samples reviewed and the quantity of annotators involved. Despite our efforts to maximize the use of available resources, these constraints may affect the scalability and reliability of our results. Additionally, we instruct all models through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2023). Exploring alternative reasoning frameworks, such as Tree of Thoughts Yao et al. (2023) or Graph of Thoughts Besta et al. (2023), could provide valuable insights into their influence on model behavior and the inferential strategies adopted. Based on our annotated data, we endeavored to develop a classifier capable of automatically identifying the inferential strategies employed in the models’ output, which was intended to complement our manual evaluation setup. However, due to the complexity of the task and limited size of our annotated dataset, our classifier struggled with generalization to new, unseen responses. In future endeavors, we aim to allocate more resources towards expanding our manual annotation efforts and explore this direction further. Finally, our study predominantly offers a behavioral analysis and does not delve into the mechanistic aspects that might explain the diversity in strategy usage by the models. Investigating how attention patterns of models with respect to specific logical connectives might influence their choice of reasoning strategy presents a compelling direction for future research.
@@ Appendix A Additional Experimental Details
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |

### v2

@@ 1 Introduction
Models are often right but for the wrong reasons: the accuracy of a model, that is the number of correct final conclusions, does not reflect whether its reasoning is sound, i.e. logically follows from the statements at hand.
@@
In contrast to humans, models occasionally adopt a symbolic strategy, where formal logical calculus is employed to solve the propositional logic problem at hand.
@@ 2 Strategies in Propositional Reasoning
To gain insights into the inferential processes humans employ in propositional reasoning, Van der Henst et al. (2002) conducted a series of experiments that study the behavior of participants during propositional reasoning. They formulated straightforward propositional logic problems with neutral content (the presence or absence of colored marbles in a box, similar to the problem illustrated in Figure 1) and requested participants to articulate their thought processes while engaging with these problems. Participants were permitted the use of paper and pencil for their workings. Both their verbal explanations and written responses were meticulously recorded, transcribed and analyzed thereafter. Van der Henst et al. (2002) discovered five strategies reasoners commonly utilize to navigate the problems, offering insights into their inferential mechanisms employed during propositional reasoning. In the following, we give a short description of each strategy (illustrated in Figure 2). For more details and additional examples, we refer to the original study by Van der Henst et al. (2002).
@@
Incremental Diagram. This strategy involves the creation of a comprehensive diagram that keeps track of all potential outcomes compatible with the premises of the problem. During the reasoning process, individuals progressively increment their diagrams to incorporate new information derived (see left box in Figure 2). The result is a single diagram that records a variety of possibilities compatible with the premises, often including even those that might be irrelevant to the task.111In contrast to Van der Henst et al. (2002), we in fact observe no single occurrence of the incremental diagram strategy in LLMs, despite the authors finding that this strategy is most frequently employed by humans. We believe that this discrepancy stems from the use of pen and paper in human assessments, implicitly encouraging diagrammatic reasoning. Exploring how this observation changes with vision-language models would be an intriguing area for future research.
@@
Supposition Following. Reasoners employing this strategy start with a supposition, e.g. by assuming a marble of a certain color. Subsequently, they trace the implications of that supposition, logically following from the premises at hand, as illustrated in the upper second box from the left of Figure 2. The result is a sequence of literals (in this case, marbles of a certain color) without logical connectives. The efficiency and success of supposition following strongly depends on the supposition made by the reasoner. While some suppositions lead to inferences that are relevant to the problem, others might lead to irrelevant conclusions.
@@
Chain Construction. When employing this strategy, reasoners construct a chain of conditional statements derived either from the premises in the problem description or from intermediate deductions. An example of chain construction is displayed in the lower second box from the left of Figure 2. Premises are converted into a chain of conditional statements that are linked by their entities. A distinctive feature of this chain is the interconnection between conditionals, where the consequent of one conditional is the antecedent of the following.
@@
Compound Strategy. Reasoners following the compound strategy combine two or more statements to derive a new compound conclusion. This process yields a series of novel conclusions, each building upon the preceding ones. An illustrative example of this strategy is given in the upper second box from the right of Figure 2. Based on the first two premises, the compound conclusion: “If blue then not red.” is inferred, and then used to draw another compound conclusion (“If blue then not red and not pink.”) together with the last premise of the problem statement.
@@ 3 Experimental Setup
Statements: 1. Either there is a white marble in the box or else there is a black marble in the box, but not both. 2. Either there is a black marble in the box or else there is a pink marble in the box, but not both. 3. There is a pink marble in the box if and only if there is a gray marble in the box. Conclusion: If there is a white marble in the box then there is a gray marble in the box.
Response:
1. First, I’ll start by analyzing the given statements:
From Statement 1, we know that the box contains either a white marble or a black marble, but not both.
From Statement 2, we know that the box contains either a black marble or a pink marble, but not both.
From Statement 3, we know that if there is a pink marble in the box, then there must be a gray marble in the box.
2. Now, let’s consider the conclusion: If there is a white marble in the box then there is a gray marble in the box.
Can we logically deduce this conclusion from the given statements?
Task Overview. Our task setup aligns with the experiment conducted by Van der Henst et al. (2002) to allow for a fair comparison between the inferential strategies found in humans and those identified in LLMs.222More specifically, experiment one of Van der Henst et al. (2002). In particular, we evaluate each model on the 12 problems of propositional logic suggested by Van der Henst et al. (2002) (an overview of each problem can be found in Figure 5 in the appendix). For each problem, models are presented with a set of statements (or premises) and must determine whether a given conclusion logically follows (for an example, see Figure 1). Eight out of 12 problems involve three premises and a conclusion, while the remaining four problems consist of four premises leading to a conclusion. All premises, as well as the conclusions resemble either biconditionals, exclusive disjunctions or conditionals. Two problems (4 and 6) include a redundant first premise. All premises are stated such that two subsequent statements contain one proposition in common, except of two problems (11 and 12), which are arranged in a non-sequential manner. For half of the problems, the conclusions logically follow from the premises, whereas for the other half, they do not. To avoid the influence of external knowledge and ensure content neutrality, Van der Henst et al. (2002) framed the problems around the presence of colored marbles in a box, with colors assigned randomly to each entity within a problem.
@@ 1. First, I’ll start by analyzing the given statements:
Evaluation Setup. We prompt each model with a system message providing context about the task they are about to solve and the format in which they should answer (for the full prompt, see Figure 5 in the appendix). Analogous to Van der Henst et al. (2002), we inform the model of its participation in an experiment designed to explore reasoning processes, and instruct it to “think aloud” as it tackles the problem. In addition to the system message, we provide a user prompt that contains the problem description. In cases where the model does not accept system messages (such as Mistral-7B-Instruct-v0.2), we prepend the content of the system message to the user prompt. To prevent biasing the model towards a certain strategy, we refrain from providing few-shot examples, as done also by Leidinger et al. (2023). Instead, we elicit reasoning through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2022). Answers for each model are generated with nucleus sampling using Llama-2-chat-hf’s default values (, temperature ), as we found this configuration to work well for all models. To account for the statistical nature of language models, we ask each model to solve the set of propositional problems across 5 random seeds, resulting in a total of 60 responses per model. Our code is publicly available at: https://github.com/mainlp/inferential-strategies.
We record all answers and manually evaluate them (a total of 300 responses) for strategies employed in their reasoning (see Figure 3 for an example). For each model response, we qualitatively evaluate for strategy and soundness. That is, we manually label the inferential strategies identified, and the logical validity of the model’s reasoning. In addition, we record whether the final answer is correct. In cases of faulty reasoning, we categorize the type of error. This comprehensive manual evaluation of model responses is independently conducted by two hired students with expertise in manual data annotation. To gauge the quality of the annotations, we report an overall Cohen’s Kappa value of . For details on the inter-annotator agreement of each label, we refer to Table 2 in the appendix. Further annotated examples can be found in Appendix C. Following the recommendations put forward by Leidinger et al. (2023), we make all input prompts, model responses and manual annotations publicly available at: huggingface.co/datasets/mainlp/inferential_strategies.
| Model | Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning |
| Zephyr-7B- | |||||||
| Mistral-7B-Instruct | |||||||
| LLaMA-2-7B | |||||||
| LLaMA-2-13B | |||||||
| LLaMA-2-70B | |||||||
| Human Reasoner† |
@@ 4 Results and Analysis
In this section, we present the results of our evaluation. We begin with a quantitative analysis of the inferential strategies employed by LLMs, as well as the logical validity of their reasoning. This is followed by a qualitative analysis providing a more in-depth examination of the models’ reasoning.
@@ 4.1 Quantitative Analysis
Table 1 provides an overview of the frequencies with which large language models employ inferential strategies when navigating the problems of propositional logic described in Section 3. Our evaluation reveals that all models display strategies akin to those observed by Van der Henst et al. (2002). In particular, we find that, similar to humans, models commonly employ supposition following, chain construction and the compound strategy. In addition, we observe that models occasionally utilize the symbolic strategy, employing techniques from logical calculus to solve the tasks (see Section 2). Note that, similar to humans, models might switch from one strategy to another during a single problem, demonstrating multiple strategies within their responses (see Figure 19 in the appendix for an example). Surprisingly, we observe that distinct model families favor different inferential strategies. For instance, Zephyr-7B- predominantly employs supposition following, while Mistral-7B-Instruct is equally inclined towards drawing compound conclusions. In contrast, models from the Llama 2 series tend to rely on supposition following and chain construction, with negligible use of the compound strategy. Our analysis further reveals a discrepancy between the correctness of the models’ final answers and the logical soundness of their reasoning. While all models achieve an answer accuracy that approximately coincides with chance in our experimental setup, an analysis of their reasoning validity reveals a different picture: LLaMA-2-70B outperforms the other models by reasoning correctly in about 31.7% of cases, while Zephyr-7B- and Mistral-7B-Instruct produce sound reasoning in 25% of the problems. We note that all models perform rather poorly on the propositional tasks, with LLaMA-2-7B failing entirely to construct sound arguments.
@@
Sound Reasoning. As previously highlighted, the accuracy of a model’s final answer does not necessarily serve as a reliable indicator of its reasoning capability. In particular, we observe that models often arrive at correct answers, but through flawed reasoning processes (refer to Figure 10 in the appendix for an illustration). Interestingly, we also find instances where models provide incorrect final answers despite reasoning correctly (for an example, see Figure 16 in the appendix). Our analysis reveals only a moderate positive correlation between the accuracy of the models’ final answers and the logical soundness of their reasoning, with a Pearson correlation coefficient and a statistically significant p-value of less than . This observation aligns with findings from previous studies Ye and Durrett (2022); Creswell and Shanahan (2022) and underscores the need for more nuanced evaluation procedures, particularly in multiple-choice settings, where models might select the correct answer by chance rather than through rigorous reasoning.
@@ 5 Related Work
Faithful Reasoning with LLMs. Large language models can be instructed to explain the reasoning process by which they derive their final conclusions Wei et al. (2022); Kojima et al. (2022). However, several studies indicate that these self-explanations might not always be faithful, i.e. accurately represent the model’s underlying reasoning process Jacovi and Goldberg (2020); Agarwal et al. (2024); Lyu et al. (2024). For instance, Turpin et al. (2023) demonstrate that LLMs such as GPT-3.5 OpenAI (2023) and Claude 1 Anthropic (2023) often fail to mention biasing features in their input that significantly influence their decisions. Instead, the models produce plausible yet misleading explanations that give a false account of the underlying decision process. Lanham et al. (2023) probe the faithfulness of explanations by evaluating how the final conclusions of LLMs change when rationales are truncated or errors are introduced. Their findings reveal that the extent to which models rely on their rationales varies strongly across models and tasks. Matton et al. (2024) propose a method to quantify the faithfulness of explanations based on high-level concepts in the models’ input that influence decision-making. By measuring the difference between the set of concepts that LLMs deem influential and the set that truly are, instances of unfaithfulness could be identified, including cases where LLMs overlook the impact of social biases in their decision-making processes. Another important consideration is whether the model’s final conclusion aligns with its preceding explanation. As highlighted in Section 4.1, a correct conclusion might not always be the product of a logically sound reasoning trace, particularly in multiple-choice setups. Conversely, a sound rationale may not always lead to a logically consistent answer. Related work by Ye and Durrett (2022) indicates that in question-answering and natural language inference tasks, explanations generated by LLMs such as OPT Zhang et al. (2022) and GPT-3 Brown et al. (2020) often do not entail the models’ final conclusions. Further studies aim to enhance the models’ faithfulness, for instance by enforcing causality from proof generation to entailment prediction. This can be achieved by either restricting the model’s context Sanyal et al. (2022); Creswell and Shanahan (2022); Radhakrishnan et al. (2023), or by utilizing deterministic tools that are inherently faithful by design Lyu et al. (2023).
@@ 7 Limitations
While our work contributes to the understanding of reasoning processes in large language models by demonstrating that these models employ inferential strategies in propositional logic similar to humans, it encompasses several limitations that could be addressed in future work.
Task setup. Our study is constrained by a limited set of problems, designed within a fixed framework that revolves around hypothesis validation based on 3-4 statements of propositional logic. We employ a constant and neutral content, disregarding potential content-effects on the models’ reasoning behavior, as shown by Dasgupta et al. (2023). Similarly, we have not yet examined factors such as the complexity of the problems, the differences between hypothesis validation and generation, and the impact of logical connectives utilized in the premises. We believe that these factors are worth investigating and leave a detailed examination to future work.
Evaluation Framework. The extent of our manual evaluation is limited by both the number of samples reviewed and the quantity of annotators involved. Despite our efforts to maximize the use of available resources, these constraints may affect the scalability and reliability of our results. Additionally, we instruct all models through zero-shot chain-of-thought prompting (“Let’s think step by step”) Kojima et al. (2022). Exploring alternative reasoning frameworks, such as Tree of Thoughts Yao et al. (2023) or Graph of Thoughts Besta et al. (2024), could provide valuable insights into their influence on model behavior and the inferential strategies adopted. Based on our annotated data, we endeavored to develop a classifier capable of automatically identifying the inferential strategies employed in the models’ output, which was intended to complement our manual evaluation setup. However, due to the complexity of the task and limited size of our annotated dataset, our classifier struggled with generalization to new, unseen responses. In future endeavors, we aim to allocate more resources towards expanding our manual annotation efforts and explore this direction further. Finally, our study predominantly offers a behavioral analysis and does not delve into the mechanistic aspects that might explain the diversity in strategy usage by the models. Investigating how model-internal mechanisms might influence their choice of reasoning strategy presents a compelling direction for future research.
@@ Appendix A Additional Experimental Details
Statements: 1. White xor black. 2. Black xor pink. 3. Pink iff gray. Conclusion: If white then gray.
Statements: 1. Brown iff orange. 2. Orange xor yellow. 3. Yellow iff green. Conclusion: If brown then green.
Statements: 1. Green iff purple. 2. If purple then gray. 3. Gray xor yellow. Conclusion: Green xor yellow.
Statements: 1. Red xor maroon. 2. Maroon xor yellow. 3. Yellow iff orange. Conclusion: If maroon then orange.
Statements: 1. Purple iff yellow. 2. Yellow iff blue. 3. Blue xor orange. Conclusion: Purple xor orange.
Statements: 1. Gray iff yellow. 2. Yellow xor olive. 3. Olive iff black. Conclusion: If yellow then black.
Statements: 1. Blue iff red. 2. Red xor white. 3. White iff pink. Conclusion: If not blue then pink.
Statements: 1. Olive xor brown. 2. Brown iff gray. 3. Gray xor maroon. Conclusion: If not olive then maroon.
Statements: 1. Purple iff blue. 2. Blue iff olive. 3. Olive xor red. 4. Red xor green. Conclusion: If purple then green.
Statements: 1. Brown iff yellow. 2. Yellow xor green. 3. Green iff purple. 4. Purple iff olive. Conclusion: If brown then olive.
Statements: 1. Red iff maroon. 2. Green xor olive. 3. Maroon iff green. 4. Olive xor brown. Conclusion: If red then brown.
Statements: 1. Blue iff brown. 2. White iff green. 3. Brown xor white. 4. Green iff purple. Conclusion: If blue then purple.
| Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning | |
| Zephyr-7B- | |||||||
| Mistral-7B-Instruct | |||||||
| LLaMA-2-7B | |||||||
| LLaMA-2-13B | |||||||